OpenStudy (anonymous):

Find the equation of the line tangent to x^5 = -xy^3 at (1,-1) Put it in slope-int. form?

OpenStudy (anonymous):

you just found the derivative right?

OpenStudy (anonymous):

think so... I can't figure out how to put it in y=mx+b form

OpenStudy (anonymous):

you need \(m\) first replace \(x\) by \(1\) and \(y\) by \(-1\) in the derivative to find \(m\) then use the point - slope formula

OpenStudy (anonymous):

So, 1^(5) = -(1)(-1)^3 ? I might not figure this one out

OpenStudy (anonymous):

no not in the equation, in the derivative you found

OpenStudy (anonymous):

im so confused.

OpenStudy (anonymous):

k lets go slow

OpenStudy (anonymous):

i can't remember the derivative exactly, but i think it was \[y'=-\frac{5x^4+y^3}{3xy^2}\] is that right?

OpenStudy (anonymous):

x^5 = -xy^3

OpenStudy (anonymous):

the derivative you found earlier, not the original equation

OpenStudy (anonymous):

hey satellite can you back when you're done here? I have a question on another problem for you

OpenStudy (anonymous):

I trust your backing mate.

OpenStudy (anonymous):

it was \[y'=-\frac{5x^4+y^3}{3xy^2}\]

OpenStudy (anonymous):

ah, yes

OpenStudy (anonymous):

now put \(x=1,y=-1\) in to the derivative

OpenStudy (anonymous):

you get \[m=-\frac{5-1}{3}=-\frac{4}{3}\] if my arithmetic is right

OpenStudy (anonymous):

that is the slope, the point is \((1,-1)\) and use the point slope formula to find the equation of the line

OpenStudy (anonymous):

so, the answer comes out to be -4/3 cause of that right?

OpenStudy (anonymous):

that is what i get for the slope, yes

OpenStudy (anonymous):

is there even a +b to the y= mx+b of it?

OpenStudy (anonymous):

It has to be formal

OpenStudy (anonymous):

you lost me there

OpenStudy (anonymous):

\[y+1=-\frac{4}{3}(x-1)\] and then put it in any form

OpenStudy (anonymous):

I have to put the eq. of the line tangent to x^5 = -xy^3 at (1,-1). Then, i gotta put it in slope-intercept form (y = mx + b)

OpenStudy (anonymous):

this is the point slope form \[y+1=-\frac{4}{3}(x-1)\] if you want it in the form \(y=mx+b\) then solve for \(y\)

OpenStudy (anonymous):

y = -4/3x + 1, right? i am terr. at math. ;)

OpenStudy (anonymous):

lol no multiply first

OpenStudy (anonymous):

y + 1 = -4/3x + 4/3x

OpenStudy (anonymous):

\[y+1=-\frac{4}{3}(x-1)\\ y+1=-\frac{4}{3}x+\frac{4}{3}\] then subtract \(1\) and get \[y=-\frac{4}{3}x+\frac{1}{3}\]

OpenStudy (anonymous):

You work miracles like no other mate. im startin' to get the gist of this

OpenStudy (anonymous):

good! yw